### v2

# Geometrical Aspects Of Resources Distribution In Quantum Random Circuits

Abstract
In this work, we explore how resources are distributed among the states generated by quantum random circuits (QRC). We focus on multipartite non-locality, but we also analyze quantum correlations by appealing to different entanglement and non-classicality measures. We compare universal vs non-universal sets of gates to gain insight into the problem of explaining quantum advantage. By comparing the results obtained with ideal (noiseless) vs noisy intermediate-scale quantum (NISQ) devices, we lay the basis of a certification protocol, which aims to quantify how robust is the resources distribution among the states that a given device can generate.
keywords:
Quantum random circuits $\cdot$ Quantum non-locality $\cdot$ Quantum entanglement $\cdot$ CHSH-Mermin-Svetlichny inequalities
[…]
[inst1]organization=University of Cagliari, addressline=Via Is Mirrions 1, city=Cagliari, postcode=09123, state=Italy
[…]
[inst2]organization=Universidad Nacional de La Plata, Instituto de Fisica (IFLP-CCT-CONICET), addressline=C.C. 727, city=La Plata, postcode=1900, state=Argentine
[…]
[inst3]organization=Technische Universität München. Institute for Advanced Study (IAS), addressline=Lichtenbergstraße 2 a, city=München, postcode=80333, state=Germany
1 Introduction
Quantum computers offer promising computational advantages for the years to come. There exist many operating devices nowadays, which can be easily accessed using the cloud quantum computing model. This rapid growth is accompanied by claims of quantum advantage with near term devices [1, 2]. Though the interpretation of the obtained results generates a heated debate and should be taken cautiously [3], it is clear that the road map is promising. There are great chances that, in the near future, quantum computers will be able to perform tasks that are very hard for classical supercomputers, even if the path to fault tolerant universal quantum computing remains challenging.
There are many hypotheses to explain why quantum computers seem to have a speed-up with regard to their classical cousins. Overall, the reasons are not completely clear. Recent studies point to different resources, such as entanglement and non-locality [4], coherence [5], and contextuality [6]. But none of these quantum features seem to suffice on its own to explain quantum advantage, due to results such as the Gottesmann-Knill theorem [7] (though see for example the discussion in [8]). Furthermore, the certification of a quantum device is usually a complicated business. As extant quantum processors are very sensitive to noise, it is important to develop different tools to assess their capabilities (see for example [9, 10, 11, 12]). The general question could be posed in this way: how are quantum resources distributed among the states that a concrete quantum device is able to produce?
Here we study the resources produced in different architectures of quantum computers using quantum random circuits (QRC) [13]. We focus on non-locality by using the Mermin [14] and the Svetlichny inequalities [15, 16]. We also study entanglement by appealing to different measures, such as tangle and concurrence (see A of this work). We first perform a theoretical study based on numerical simulations using Qiskit [17] and Amazon Braket SDK [18]. We analyze the performance of the devices with different levels of noise and number of shots, and compare universal (Clifford $+$ T) vs non-universal (Clifford) sets of elementary gates. Given a set of physical features which are thought of as a resource for quantum computing (as for example, non-locality), we quantify to which degree a device can generate a state space rich enough to generate the desired resources.
The paper is organized as follows. In Section 2 after giving a very brief overview of QRC, we show the results of the simulations of violation of Mermin and Svetlichny inequalities for different qubits numbers. In Section 3, we present results regarding entanglement for comparison purposes. Finally, in Section 6, we draw our conclusions.
2 Quantum Random Circuits: distribution of resources
QRC are used in different areas of quantum information science [13]. In order to give a general idea, consider a scheme of circuit generation as depicted in Figure 1. On each stage of the protocol, random unitary gates are chosen and applied to the circuit. As a result, a random circuit is obtained. Remarkably, QRC are a promising tool to show quantum advantage [1]. Here, we use them to show how quantum resources are distributed among the possible states that a quantum computer is able to generate. An ideal quantum computer should be able to reach all possible states of a system of $N$-qubits. A real quantum device is able to implement a given set of elementary quantum gates on a fixed set of qubits. By generating QRC with different depths, sets of elementary gates, and levels of noise, we study how robust is the resources distribution with regard to the ideal case.
Specifically, by comparing Clifford (H, S and CNOT) vs non Clifford (Clifford $+$ T) sets of gates, we highlight the differences of how quantum resources are distributed between universal vs non-universal quantum computing.
Figure 1: We used the Amazon Braket SDK [18] to develop a function that generates QRC. Schematic representation of the random choice of gates. We first fix a set of elementary gates, and then, on each layer of the circuit apply them randomly over the different qubits. In the schematic diagram, we display different types of one-qubit gates (boxes indicating H, T and S gates), together with two-qubits entangling gates (being CNOT the paradigmatic example). After $N$ layers are applied, the output statistics for different observables is drawn, and the degree of violation of CHSH, Mermin and Svetlichny inequalities can be checked.
In this section we focus on non-locality, as is expressed by the violation of certain inequalities. In the following section, we will deal with entanglement.
We start by analyzing a simple two-qubits system and the violation of CHSH inequalities [19]. For the state associated to each generated QRC, we compute its maximal degree of the violation. We consider from now on an experimental situation in which two dichotomic observables $A_{1}$, $A_{1}^{\prime}$, $A_{2}$ and $A_{2}^{\prime}$, with outputs $1$ and $-1$ each, can be performed on each system. Let
$S_{2}=E\left(A_{1}A_{2}\right)-E\left(A_{1}A_{2}^{\prime}\right)+E\left(A_{1}^{\prime}A_{2}\right)+E\left(A_{1}^{\prime}A_{2}^{\prime}\right),$ (1)
where $E(A_{1}A_{2})$, for example, is the correlation function between $A_{1}$ and $A_{2}$, and so on with the primed variables. On the basis of the so called local realism (or local hidden variable model) it is possible to obtain the CHSH inequality [19]:
$|S_{2}|\leq 2.$ (2)
Delving into the realm of quantum mechanics, our focus on dichotomic measurements allows us to confine our exploration to two-dimensional systems. In this case the operator inside the modulus of the left-hand side of Eq.2 becomes:
$O_{2}=\vec{\sigma}\cdot\vec{a}_{0}\otimes\vec{\sigma}\cdot\vec{b}_{0}-\vec{\sigma}\cdot\vec{a}_{1}\otimes\vec{\sigma}\cdot\vec{b}_{0}+\vec{\sigma}\cdot\vec{a}_{0}\otimes\vec{\sigma}\cdot\vec{b}_{1}+\vec{\sigma}\cdot\vec{a}_{1}\otimes\vec{\sigma}\cdot\vec{b}_{1},$ (3)
where $\vec{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ with $\sigma_{x}$, $\sigma_{y}$ and $\sigma_{z}$ that are the Pauli operators and $\vec{a}_{0}$, $\vec{a}_{1}$, $\vec{b}_{0}$ and $\vec{b}_{1}$, are unit vectors representing the orientations of the polarizers of each pair. To each direction there is a pair of angles associated. For a given state $\rho$ we maximize the quantity
$V_{2}(\rho)=|\mbox{Tr}(\rho O_{2})|,$ (4)
among possible values of the angles. If the local realism is violated and consequentely non locality is present in a given state $\rho$ with regard to the chosen family of observables [^1] Notice that, in principle, there could exist a more general family of observables for which a non-locality inequality is violated. We are restricting here the analysis to the CHSH type observables (with variable angles)., there should exist a combination of angles for which $S_{2}(\rho)=V_{2}(\rho)>2$.
Now we turn to the study of systems with 2, 3, 4 and 5 qubits. When more than two particles are present, there exist different types of non-local states. First, one could start asking the question whether there exists a local hidden-variable model reproducing the correlations of the system under study. This problem was addressed by Mermin [14] (see also [20] and [21]). By requiring the condition of local hidden variable model, it is possible to derive a set of inequalities based on the so-called Mermin polynomials $M_{N}$ (see for example [20], section II-C). These polynomials are recursively defined in the following way
$M_{2}=\frac{1}{2}\left(A_{1}A_{2}+A_{1}^{\prime}A_{2}+A_{1}A_{2}^{\prime}-A_{1}^{\prime}A_{2}^{\prime}\right),$ (5)
and
$M_{N}=\frac{1}{2}M_{N-1}\left(A_{N}+A_{N}^{\prime}\right)+\frac{1}{2}{M}_{N-1}^{\prime}\left(A_{N}-A_{N}^{\prime}\right),$ (6)
where $M^{\prime}_{N-1}$ is obtained from $M_{N-1}$ by exchanging all the zero-indexed and and one-indexed $A$’s. The local realism limit is given by [20]:
$|\langle M_{N}\rangle|\leq 1.$ (7)
For $N=2$ we recover one of the equivalent ways of writing the CHSH inequalities up to factor $1/2$. The violation of the $N$-particle Mermin inequality implies that the correlations cannot be modeled using local realism. But, in a multi-partite system, restricting the analysis to the simple absence of local hidden-variable models might yield a narrow perspective on the possible correlations involved. For example, one could have a state of three particles in which the first two are maximally correlated, while there is no correlation with regard to the third one, as is the case for a quantum system of three particles prepared in the state
$|\psi\rangle=\frac{1}{\sqrt{2}}\left(|00\rangle+|11\rangle\right)|0\rangle.$ (8)
The above state clearly displays non-local correlations between the first and second particles, but these correlations do not involve all the parties. In order to distinguish among the different types of correlations, G. Svelitchny and collaborators [15, 16] developed a set of inequalities which, if violated, will reveal the presence of genuine multipartite non-locality [^2] See also the discussion presented in Ref. [23].. The Svetlichny inequalities have also been experimentally confirmed (see for example [22]).
Consider the correlation functions $E(A_{1}^{(i)}A_{2}^{(j)}A_{3}^{(k)})$, which represent the expected value of the product of the measurement outcomes of single particle random variables $A_{1}^{(i)}$, $A_{2}^{(j)}$ and $A_{3}^{(k)}$ ($i,j,k$ are primed or not primed variables). Define
$S_{3}$ $=E\left(A_{1}A_{2}A_{3}\right)+E\left(A_{1}A_{2}A_{3}^{\prime}\right)+E\left(A_{1}A_{2}^{\prime}A_{3}\right)+E\left(A_{1}^{\prime}A_{2}A_{3}\right)$ (9)
$-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}^{\prime}\right)-E\left(A_{1}^{\prime}A_{2}^{\prime}A_{3}\right)-E\left(A_{1}^{\prime}A_{2}A_{3}^{\prime}\right)-E\left(A_{1}A_{2}^{\prime}A_{3}^{\prime}\right),$
then, the three-qubits Svetlichny inequality reads:
$|S_{3}|\leq 4.$
If a state violates the above inequality, then, it presents genuine three-partite non-locality.
In a quantum setting, consider the emission of triplets of particles in a pure quantum state from a source, for which their state is possibly unknown. During each execution of the experiment, one of the two possible alternative measurements is performed: $A_{1}$ or $A_{1}^{\prime}$ on the first particle, $A_{2}$ or $A_{2}^{\prime}$ on the second particle, and $A_{3}$ or $A_{3}^{\prime}$ on the third particle. Each operator is of the form $\vec{\sigma}\cdot\vec{a}_{i}$, $\vec{\sigma}\cdot\vec{b}_{i}$ and $\vec{\sigma}\cdot\vec{c}_{i}$, with $\vec{a}_{i}$, $\vec{b}_{i}$ and $\vec{c}_{i}$, different unit vectors (with $i=0,1$). For the quantum case, we have $E(A_{1}A_{2}A_{3})=\mbox{tr}(\rho A_{1}\otimes A_{2}\otimes A_{3})$, with $A_{1}$, $A_{2}$ and $A_{3}$ Hermitian operators. Define the three-qubits Svetlichny operator by:
$O_{3}$ $=A_{1}\otimes A_{2}\otimes A_{3}+A_{1}\otimes A_{2}\otimes A_{3}^{\prime}+A_{1}\otimes A_{2}^{\prime}\otimes A_{3}+A_{1}^{\prime}\otimes A_{2}\otimes A_{3}$ (10)
$-A_{1}^{\prime}\otimes A_{2}^{\prime}\otimes A_{3}^{\prime}-A_{1}^{\prime}\otimes A_{2}^{\prime}\otimes A_{3}-A_{1}^{\prime}\otimes A_{2}\otimes A_{3}^{\prime}-A_{1}\otimes A_{2}^{\prime}\otimes A_{3}^{\prime}.$
For a given three-qubits state $\rho$, we maximize the quantity
$V_{3}(\rho)=|\mbox{Tr}(\rho O_{3})|,$ (11)
for all possible values of the polarizers. As in the two-qubits case, a violation of the above inequality implies that the state has genuine three-partite non-locality. Thus, given a state $\rho$, we look for the angles which maximize (11), and use that maximal value to quantify the eventual non-locality present in the state. We refer to the value obtained as violation level. A state with no non-locality with regard to the chosen set of observables, will display a violation level below the classical limit of 4. As is well known [15], for the three-qubits case, the maximum value of violation that can be reached with a quantum state is $4\sqrt{2}$. Thus, it is the maximum violation level reachable using quantum states.
For the case of an arbitrary number $N$ of qubit systems, the inequality is built using the quantity [16]:
$S_{N}^{\pm}=\sum_{I}\nu_{t\left(I\right)}^{\pm}A^{i_{1}}_{1}\cdots A^{i_{N}}_{N},$ (12)
In the above inequality, $\nu_{t}^{\pm}=\left(-1\right)^{\frac{t\left(t\pm 1\right)}{2}}$, with $I=\left(i_{1},i_{2},\cdots,i_{N}\right)$, with each $i_{j}$ indicating whether a prime appears or not in the corresponding variable. The number $t\left(I\right)$ indicates the number of times primes appear in I, and $\nu_{t}^{\pm}$ is a sequence of signs. Two types of inequalities are derived, namely
$|S_{N}^{\pm}|\leq 2^{N-1}.$ (13)
We now show the numerical results. In Figures 2 and 3 we show histograms of the obtained violation levels of the Svetlichny and Mermin inequalities, respectively, for states generated using QRC from two to five qubits [^3] The histogram for the CHSH inequality is repeated in both Figures, in order to make the comparison easier for the reader.. The histograms are obtained for $100.000$ randomly generated sates. We do this for the non-universal set formed by Clifford gates, a universal set, adding T to the previous set, and random unitaries generated using the python library SciPy. The latter simulate the generation of random unitaries distributed according to the Haar measure [^4] The Haar distribution is a translation-invariant probability measure. Intuitively speaking, it is a uniform probability distribution in the space of unitary matrices.. In Figure 4 we show the violation levels obtained for different levels of noise for a two-qubits system and the CHSH inequality, and for a three-qubits system with respect to the Mermin and Svetlichny inequalities.
It is important to remark that, in this work, the states generated using the Clifford set have a very different geometrical structure with regard to those generated with Clifford $+$ T. While in the latter the states generated tend to cover the entire set of quantum states (and it is very unlikely to obtain a given state twice), in the former, it is indeed likely to obtain repeated states (when the circuits are chosen at random). Thus, for the Clifford set, the term “random” only applies to the way in which the gates of the elementary set are chosen to form each circuit (i.e., following the procedure indicated in Figure 1), and should not be confused with that of uniform distribution which applies when using random unitaries (and which the universal set resembles). All possible states generated using Clifford gates can be actually computed, and the sets of states having the different possible values of degree of violation have the same weight as in the histograms displayed here. The states associated to the Clifford set display a rich and intriguing geometrical structure that we will study in a separate work.
Figure 2: Violation levels of the Mermin inequality obtained for 100,000 random circuits: (a) two-qubits, (b) three- qubits, (c) four-qubits and (d) five-qubits using Clifford gates, Clifford $+$ T gates, and random unitaries following the Haar distribution.
Figure 3: Violation levels of the Svetlichny inequality obtained for 100,000 random circuits: (a) two-qubits, (b) three-qubits, (c) four-qubits and (d) five-qubits using Clifford Gates, Clifford $+$ T gates, and random unitaries following the Haar distribution.
Figure 4: Histograms of violation level for $3.000$ random circuits of two and three-qubit systems with a depth of 41 layers and a depolarizing channel with parameters 0.00002, 0.00009, 0.0003, 0.0009 and 0.001, for one qubit gates, respectively. In (A) we show the two-qubits case using the CHSH inequality. For the three-qubits systems we show the histograms for (B) Mermin inequality and (C) Svetlichny inequality.
The plots of Figure 4 clearly indicate how the resources distribution depend on the noise level. The comparison between universal vs non-universal sets of gates indicates that the main difference comes from the way in which resources are distributed. In both cases we obtain the maximum possible value of violation, which is attained by the GHZ state. This means that even a quantum device which is classically simulable can reach maximal non-locality. But, in a universal device, the different degrees of non-locality are homogeneously distributed among all possible values available in the full quantum state space. This is in line with previous works (compare with [24]). By augmenting noise, one sees that the ideal histograms are deteriorated in both cases [^5] Notice that when noise is increased the histograms tend to a Gaussian-like distribution supported by non-violating values.. By comparing how an empirical histogram departs from the ideal case, one could certify how robust is the resources distribution in a concrete quantum device. The above analysis can be considered then as the basis of a certification protocol.
It is interesting to look at the violation fraction [^6] The violation fraction refers to the percentage of the states generated which violate a given inequality for some set of angles. as the number of qubits is increased (see Figure 5). For the Svetlichny inequality, we find that the violation fraction tends to zero as the number of qubits grows. This indicates that the relative volume of the set of genuine multipartite non-local states becomes very small as $N$ grows. In other words, the probability that a randomly picked pure state displays genuine non-local multipartite correlations, becomes very small in that limit. Contrarily, for the Mermin inequality, the violation percentage tends to remain very high as $N$ grows, indicating that the volume of the set of states displaying some form of non-local correlation is big (around $90\%$ of the whole state space). This is in agreement with the fact that the volume of the set of separable states tends to zero as $N$ goes to infinity. Notice that, given that entanglement does not necessarily entail non-locality, the results presented here are slightly different from those related to entanglement.
Figure 5: Violation fraction for Mermin and Svetlichny inequality as a function of the number of qubits using Clifford gates, Clifford $+$ T gates, and random unitaries following the Haar distribution.
In particular, our results indicate that the volume of the set of states that display no non-local correlation (at least with regard to the observables appearing in Mermin’s inequality) does not seem to tend to zero as $N$ grows. Here we are offering numerical evidence, but our findings give place to the question of how to prove theoretically that this is so. We will address all these questions in future works.
It is also important to take into account the circuits depth. In Figure 6 we show the violation levels obtained for Mermin and Svetlichny inequalities from two to five qubits systems varying the circuit’s depths (without noise). For a higher depth, we find curves that resemble those of random unitaries. But for lower depth levels, there are more chances of obtaining states with higher maximal violation levels (when the gates are randomly picked). This dependence could be of practical use when analysing concrete devices.
Figure 7 gives a concrete snapshot of how the system behaves with regard to different levels of noise and depth. As expected, as noise grows, the violation fraction becomes smaller —until eventually, it becomes zero. Notice also that, for a given noise level, there is a concrete depth for which the violation fraction becomes maximal. This is due to the fact that noise accumulates each time a quantum gate is applied. A low depth will give place to a poor family of states but, if the depth is too high, noise starts to play a significant role, and destroys all resources.
Figure 6: Representation of noiseless violation level histograms depicting the cases of (A) CHSH inequality, (B) Mermin inequality and (C) Svetlichny inequality across different system sizes, including two-qubit, three-qubit, four-qubit, and five-qubit configurations, as well as various gate depths. The scope of gate operations is confined to the Clifford $+$ T set. Notice that for higher depth levels, the curves resemble those of states generated with random unitaries. But, in this case, the chance of obtaining a state with a near maximal violation level tends to decrease as the number of qubits grow. On the contrary, for lower depth levels, it becomes more likely to obtain such states.
It is also interesting to look at the violation levels for a higher number of qubits and the Clifford set only. The results are displayed in Figure 8. Using the Clifford set, it is easier to achieve maximal violation values with very few iterations. Again, we recover a pattern where the violation levels obtained for the different states are concentrated in specific values. For a better visualization of this fact, we have plotted the list of values obtained against themselves for the different numbers of qubits in Figure 8. The violation levels concentration is the most salient difference between universal and non-universal sets of gates. Remarkably, the number of pick values obtained seems to be proportional to the number of qubits.
Figure 7: (A) CHSH inequality, (B) Mermin inequality and (C) Svetlichny inequality are examined, wherein figures (a) and (b) delineate the violation level and the fraction of violations as functions of circuit depth for Clifford gates, across varying noise levels. Figures (c) and (d) present analogous analyses, albeit for the Clifford $+$ T gate set.
Figure 8: (a) Histograms of violation levels (without noise) of Mermin inequality for 100,000 random circuits for $N=2,3,4,5$ for states generated using the Clifford group. In (b) we plot the list of violation levels obtained for the Mermin inequality against themselves for $N=2,3,4,5$. (c) and (d) the same as in the previous case but for Svetlichny inequality. If all possible values were reached, then, we should observe a line segment. But, as remarked above, the figures obtained for the Clifford group have “holes", in the sense that the violation levels tend to concentrate around specific points. This is related to the fact that the states generated are concentrated in turn in specific regions of the quantum state space. These plots should be compared with those of the histograms in Figures 2 and 3 (and with the results presented in [24] and [25]).
3 Entanglement measures and other quantifiers
Up to now, we have focused in non-locality, quantified as the violation levels of Mermin and Svetlichny inequalities. For the sake of comparison, it is also interesting to analyze what happens with entanglement. A short overview of the entanglement measures used in this work can be found in A. In Figure 9 we show how entanglement is distributed for different different entanglement measures and sets of elementary gates for three-qubits QRC.
In Figure 12, we plot the values of entanglement obtained for different measures and sets of gates, against the violation levels of the three-particle Svetlichny inequality. This gives an idea about how entanglement and non-locality are related. As expected, in both figures, the results obtained for the complete set are clearly equivalent to those obtained by ideal random unitaries distributed according to the Haar measure. Also, both graphics clearly illustrate the difference between complete vs incomplete sets of quantum gates. We recover again the following result: in non-universal sets of gates, the resources (in this case, entanglement) are highly concentrated in certain particular values. For universal sets, on the other hand, a vast plurality of values is obtained, indicating that, using QRC, the full potentiality of the quantum state space is obtained.
Figure 9: 50,000 random circuits for a three-qubits system with a depth of 250 layers. Distribution for different types of entanglement applied on three different groups of quantum circuits: Clifford, Clifford $+$ T and unitary random with Haar measure.
In Figure 11, we plot the violation levels of Mermin and Svetlichny vs the tangle of $50.000$ three-qubit states generated using quantum random circuits with a universal set of gates. These plots extend and can be compared with the results presented in previous works [26].
Finally, in Figure 11, we show the behavior of quantum magic based on the stabilizer $\alpha$-Rényi entropy. Roughly speaking, this quantity describes how many non-Clifford gates are necessary to generate a given quantum state. As the number of qubits is increased, the histograms seem to become narrower and more picked.
Figure 10: 50.000 three-qubit circuits using noiseless QRC with a universal set of gates. The red dashed lines depicted in the graphs correspond the family of states studied in Ref. [26] (we include them here for comparative analysis).
Figure 11: Distribution of Quantum Magic across 100,000 randomly generated circuits for three, four, and five-qubit systems, featuring random depths. The circuits were constructed utilizing gates from the Clifford $+$ T set and random unitaries, distributed in accordance with the Haar measure. Notably, for circuits constituted exclusively of the Clifford group, the quantum magic value is quantified as zero, hence its omission from the display. It is important to stress here the relevance of the depth of the circuits generated. One could fix a reasonable depth level chosen by inspection. Also, one can set the depth as a random parameter when generating the circuits, as we have done in the above plots. The latter option yields results which resemble those of random unitaries with more precision.
Figure 12: The various metrics of quantum correlations of the generated states are graphically represented along with the violation levels of the Svetlichny inequalities for three-qubit systems across 50,000 randomly generated circuits. The red dashed lines featured in the graphs correspond to the family of states studied in references [27] and [28] (included here to facilitate a comparative analysis).
4 Certifying quantum processors
The results obtained in the previous sections indicate that there is a sharp and quantifiable difference between, first, universal vs non-universal sets of gates, and second, ideal (noiseless) and non-ideal machines. By comparing the ideal vs experimentally obtained histograms of violation degrees, one can have a global measure of to what degree a quantum processor is able to generate the whole quantum state space. This suggests a methodology for testing the performance of a device, following the steps below:
  1. 1.
    Input: Specify the number $N$ of quantum random circuits generated, number of qubits to be used $Q$, depth of the circuit $D$ and number of shots $S$ to be used.
  2. 2.
    Implementation on the Device Using the native gates of the device, generate $N$ quantum random circuits ($S$ shots per circuit) and perform the measurements needed for the chosen resource to be tested (for the case, of a Bell-type inequality, we should perform the measurements related to the operator associated to that observable).
  3. 3.
    Comparison with ideal device: Compare the obtained results with the ideal ones (using the same set of gates). It could be to perform a histogram, compute the probabilities associated, and use some entropic measure to compare.
Implementation on quantum processors
Here we show an example of implementation in IonQ: Aria-1 and IQM: Garnet processors. Since implementing thousands of states in a real quantum processor would be expensive, here we focus on a variant of the steps mentioned in Methods and restrict ourselves to three-qubit systems. In the third step, instead of implementing all generated states, we proceed as follows. First, we divide the interval $[4,5.6]$ (the violation values range) in subintervals of length $0.2$. Next, from each interval $I$, we look at the set $\Delta_{I}$ of generated states that have a maximal violation value in the interval $I$. From the set $\Delta_{I}$, we pick a representative state minimizing the number of gates and the number of entangling gates (CZ in the IQM: Garnet and MS in the IonQ: Aria-1 case).
Figure 13: Schematic representation of IQM Garnet’s connectivity and the spin orientations measured in one of our experiments. Each arrow represents a possible spin direction of qubits $1$, $2$, and $4$. All possibilities are listed in the right.
In order to generate the random circuits, we modified the function used in the previous section to generate circuits of variable length, and connectivity of each processor and using only native gates (PRX and CZ for IQM: Garnet, and GPi, GPi2 and MS for IonQ: Aria-1). Local spin observables are implemented as proper rotations with PRX or GPi-GPi2 gates depending on the processor (the theoretical angles obtained for the synthetic data can be easily translated in terms of those gates). A schematic picture of the IQM: Garnet connectivity and the different spin directions used in a test of a Svetlichny violation is shown in Figure 13. The results of running ten states are displayed in Figure 15, together with their corresponding theoretical values of maximal violations.
It is also instructive to see what happens when the same violation experiment is repeated several times. In Figure 15 we show a plot with ten repetitions for each state of Figure 15. The results obtained indicate that while a violation of the Svetlichny inequality is clearly obtained, the results are far from the ideal case (compare these results with those of Figure 3).
Figure 14: Comparison between IonQ: Aria-1 (green dots) and IQM: Garnet (red dots). For each interval, we chose a representative state which uses a minimal number of MS and CZ gates respectively (and a minimum global number of gates). For each of them, we store the values of angles of spin measurement of (theoretically) maximal violation, in which one instance of experiment is chosen for each QPU.
Figure 15: Degree of Svetlichny violation for different instances of three qubits random circuits implemented in IQM Garnet. We have prepared ten repetitions of each state to study the fluctuation in the different runs of the experiment.
As the results of each instance of a Svetlichny violation fluctuate, we have performed an analysis of the dependence with the number of shots. The results are displayed in Figure 16. Clearly, the fluctuation tends to decrease as the number of shots grow. But using $1000$ shots gives a good idea of the degree of violation attained by the QPU in the mean.
Before we go on, it is important to make a digression about the interpretation of the results presented in Figure 15. When trying to generate a state $\rho$ in a quantum computer, one needs to implement a series of instructions $I=\{i_{1},i_{2},....,i_{n}\}$ in the quantum processing unit.
Figure 16: We have prepared ten repetitions of each state.
Let us call the state resulting from those instructions $\rho^{ideal}_{I}$. If there were no imperfections present during the process, then, we should have $\rho=\rho^{ideal}_{I}$. But in an actual experiment, $\rho$ and $\rho^{ideal}_{I}$ might be significantly different. Therefore, the implementation of the instructions is automated in such a way to optimize the task of creating a state as similar as possible to $\rho$. For that reason, the actual instructions implemented in the QPU will be denoted by a list $I^{\prime}=\{i_{1}^{\prime},i_{2}^{\prime},....,i_{n^{\prime}}^{\prime}\}$. While $I^{\prime}$ might be quite different from $I$, if the optimization process is successful, the resulting state $\rho^{real}_{I^{\prime}}$ will be very similar to $\rho$. In other words, we end up with $\rho^{real}_{I^{\prime}}\approx\rho^{ideal}_{I}$. Notice that the notation $\rho^{real}_{I^{\prime}}$ refers to the actual state generated in the QPU (which might be in turn different from the ideal state $\rho^{ideal}_{I^{\prime}}$ produced with noiseless instructions $I^{\prime}$). The situation is similar when we implement measurements. Suppose that in order to emulate a certain measurement in a QPU we need to implement the list of instructions $M=\{m_{1},m_{2},...,m_{l}\}$. Let us call the resulting ideal state $\rho^{ideal}_{I+M}$ (i.e., the state that results from implementing noiseless instructions). Again, in an actual experiment, we would obtain $\rho^{real}_{I^{\prime}+M^{\prime}}$. Ideally, to perform the desired measurement in the state $\rho$, we should run the circuit with the instructions $I\cup M$ and sample in the computational basis. As performing a measurement in a basis different than the computational one is equivalent to performing a suitably chosen rotation and measuring in the computational basis, if the optimization process works reasonably well, one concludes that $\rho^{real}_{I^{\prime}+M^{\prime}}$ is the best representative of the ideal process of preparing $\rho$ and measuring $M$. Notice also that in two different runs of the instructions $I\cup M$, one might obtain different circuits and, in principle, the chosen qubits might be even different. One could say that this is not relevant, because the goal is to reproduce the different copies of $\rho^{real}_{I+M}$ with the highest possible fidelity.
Figure 17: Degree of violation for GHZ states in Aria-1 and Garnet for two to five qubits. The statistics for the correlations of each shot are computed with 1000 shots.
It is important to notice that the above discussion can be applied mutatis mutandis to every quantum experiment, since every quantum experiment actually performed can be reduced to a list of operations on a physical system. The quantum systems used might be different (for example, different instances of photons or atoms), and the implementation of the procedures and measurements might slightly differ from one repetition to another. But the main goal of every experiment is to generate several copies of the ideal state and measurements (which are as similar as possible to it), in order to perform the required statistics.
Using IonQ’s, IQM’s transpilation system and IQM: FakeAdonis simulator we performed a final experiment by sending GHZ states. The results are presented in Figure 17.
5 Genuine multipartite non-locality vs contextuality
The experimental test of Bell inequalities was challenging due to the presence of possible loopholes: defects in the experimental implementation can be used to give explanations of the observed data in which the local-realism hypothesis can be maintained. After several years of huge efforts, the so-called loophole-free Bell tests were finally announced [29, 30, 31]. Taking into account how hard it was to reach that experimental degree of certification, one may wonder up to which extent it is realistic to affirm that genuine multipartite non-locality was produced in a given quantum processor. Indeed, in NISQ devices, qubits are usually very close, and the measurement of the different parties seem to be far from the conditions required for a loophole-free non-locality test. A quick thought indicates that it is unlikely that anyone can construct in the near future a quantum processor in which genuine multipartite non-locality can be tested in a loophole-free way. Therefore, we should be conservative about the certification of non-locality in current quantum processors (see also the discussion in [32], page 2). Still, the violation of Svetlichny inequalities in quantum processors can be considered as a form of non-classicality, given that it discards a certain form of hidden variable models. Perhaps, it is safer to say that they represent a particular form of contextuality in the following sense. For a certain family of observables (the local spin orientations with suitably chosen angles), there exists no global probability distribution satisfying certain constrains. Alternatively, one possible strategy is to work under the no-cross-talk assumption [32]. In any case, it is reasonable to consider the violations presented in the previous section as a form of non-classicality, given that they imply that the state presents genuine multipartite entanglement.
The above observations give place to a very interesting debate about the foundations of quantum physics that has technological implications. A proponent of the thesis that loophole-free certified non-locality is a necessary and sufficient condition for quantum advantage, will have to face the problem that it is extremely unlikely that a quantum processor will be ever able to give it for granted. This has implications on the development strategy of QPUs since, in the quest for quantum advantage, it would be necessary to move forward under the assumption that non-locality is actually present. Otherwise, efforts should be devoted to testing this feature in a loophole-free way, but this doesn’t seem realistic with the extant technologies and experimental techniques. On the contrary, a defender of the thesis that quantum advantage is all about contextuality will still have a point in moving forward. One could always consider a quantum processor as a simulator of genuine multipartite non-locality that uses contextuality as a resource. Is this enough for quantum advantage? Only time can tell the answer to this intriguing question. We will address this fascinating subject with more detail in a future work. With this short section, we just want to draw the reader’s attention on the relevance of this problem, since it connects both, foundations of quantum physics debates, and technological developments. We think that the results presented in this work contribute to a better understanding of the challenges that the development of quantum processors must face.
6 Conclusions
In this work we have characterized the differences between the Clifford (Classically simulable) and Clifford $+$ T gates with regard to the violation of the Svetlichny and Mermin inequalities. Interestingly, even with a non-universal gate set, the maximal possible value of violation is obtained. Also, in the presence of noise, the Clifford set is still able to produce near maximal violations, as expected. The remarkable fact is that the main difference between the universal and the non-universal set of gates is that the latter fails to cover all possible values of the resource available in the quantum state space. This was clearly shown in Figure 8, with the presence of “holes” in the list of possible values of the resource. One could say that, essentially, the difference between universal vs non-universal quantum computing resides in the fact that, geometrically, the resources are distributed very differently along the quantum state space.
Another relevant result is that the (relative) volume of the set of states violating the Svetlichny inequality tends to decrease substantially as the number of qubits increases. In fact, it becomes almost zero already for five qubits (see Figure 5). This is in contrast with the asymptotic behavior of the volume of the set of entangled states —quantifying entanglement with the von Neumann entropy. As is well known, the volume of the set of separable states tends to zero as $N$ increases. Also, we have found that the violation fraction for Mermin’s inequality is still non-negligible for five qubits. It is important to have in mind that the fact that even if a given state does not violate Svetlichny’s inequality, it is not granted that there will not exist a stronger inequality revealing genuine multipartite non-locality for that state. Our findings open the door to keep inquiring on this subtle topic, indicating a complex behavior of the geometry of the set of quantum states as the number of qubits increases.
By examining how non-locality is distributed in QRC for different gates sets, we have also gained insights on the impact of noise on near term quantum devices. We also analyzed the dependence of the results obtained with regard to circuits depth, and compared with different measures of entanglement and quantum magic. Regarding entanglement measures, our results are consistent with previous efforts, extending the scope of the analysis to states which are uniformly distributed along the quantum state space. Globally, our results go beyond previous works in characterizing the connections between entanglement and non-locality.
Our findings could serve as the basis for developing certification protocols in future works, as they provide valuable information on the quantum computer’s ability to produce relevant resources —such as genuine multipartite non-locality.
CODE AVAILABILITY
The code used to simulate the QRC is available at https://github.com/QuantumUnica/QRC_Search
Appendix A A brief summary of the measures used in this work
In this section we give a brief exposition of the entanglement measured used in this work.
A.1 Entanglement entropy
Given a quantum state $\rho$, its Von Neummann entropy is defined as
$S\left(\rho\right)=-\sum_{i}\lambda_{i}\ln\lambda_{i},$ (14)
where $\lambda_{i}\geq 0$ are the eigenvalues of $\rho$. The entanglement entropy of $\rho$ is computed as the von Neuman entropy of the reduced state of the system.
A.2 Negativity
It is possible to detect the presence of entanglement between subsystems using the negativity $\mathcal{N}$. In this case where we focus on tripartite systems, it is known as tripartite negativity, and is defined as follows [33]
$\mathcal{N}\left(\rho^{tC}\right)=\sum_{i}\left|\lambda_{i}\left(\rho^{tC}\right)\right|-1,$ (15)
where $\rho^{tC}$ is the partial transpose of $\rho$ with respect to the subsystem C, and $\lambda_{i}\left(\rho^{tC}\right)$ are the eigenvalues of $\rho^{tC}$. The negativity can be equivalently interpreted [34] as the sum of the absolute values of the negative eigenvalues of $\rho^{tC}$.
A.3 Concurrence
The multiqubit concurrence [35] for a pure three-qubit state $|\psi\rangle$ is given by
$C_{3}\left(|\psi\rangle\right)=\sqrt{1-\frac{1}{3}\sum_{i=1}^{3}\mbox{Tr}\left(\rho_{i}^{2}\right)}.$ (16)
A.4 Multipartite systems with bipartite entanglement
As computing the Von Neumann entropy can pose computational challenges due to the need to compute the eigenvalues of the reduced density operator, it is often preferable, particularly in the context of multipartite systems, to utilize an entanglement measure derived from the linear entropy:
$Q=2-\left(\frac{2}{n_{q}}\right)\sum_{i=1}^{n_{q}}\mathrm{Tr}[\rho_{i}^{2}],$ (17)
here, the $\rho_{i}=\mbox{Tr}_{i}|\psi\rangle\langle\psi|$ represent the reduced density matrix of the $i$-th qubit, obtained by tracing out the remaining $n_{q}-1$ qubits. The parameter $n_{q}$ denotes the total number of qubits. It follows that $0\leq Q\leq 1$. The distribution of $Q$ generated by a set of random unitary states, along with Haar measurements, is depicted in Fig. 18.
Figure 18: Entanglement distribution for the states generated with our QRC function for eight qubits and different circuit depths. The image on the inset was taken from [12] and is included here for comparison.
A.5 Tangle
The 3-tangle (or residual tangle) coincides with the modulus of a so-called hyperdeterminant which was introduced by Cayley [36]
$\tau_{3}=4\left|H\det\left(t_{ijk}\right)\right|,$ (18)
For our work, the hyperdeterminant $H\det\left(t_{ijk}\right)$ is a polynomial of order four in the amplitudes $t_{ijk}$. It can be expressed by using the wave function coefficients $\left\{\psi_{000},\psi_{001},\cdots,\psi_{111}\right\}$ [37] as
$\tau_{3}$ $=4\left|d_{1}-2d_{2}+4d_{3}\right|$
$d_{1}$ $=\psi_{000}^{2}\psi_{111}^{2}+\psi_{001}^{2}\psi_{101}^{2}+\psi_{100}^{2}\psi_{011}^{2}$
$d_{2}$ $=\psi_{000}\psi_{111}\psi_{011}\psi_{100}+\psi_{000}\psi_{111}\psi_{101}\psi_{010}+\psi_{000}\psi_{111}\psi_{110}\psi_{001}$
$+\psi_{011}\psi_{100}\psi_{101}\psi_{010}+\psi_{011}\psi_{100}\psi_{110}\psi_{001}+\psi_{101}\psi_{010}\psi_{110}\psi_{001}$
$d_{3}$ $=\psi_{000}\psi_{110}\psi_{101}\psi_{011}+\psi_{111}\psi_{001}\psi_{010}\psi_{100}.$
A.6 Genuine measurement entanglement with concurrence
For $n$-partite pure states $|\Psi\rangle\in H_{1}\otimes H_{2}\otimes\cdots\otimes H_{n}$, on [38] define the gme-concurrence as
$C_{gme}\left(|\Psi\rangle\right):=\min_{\gamma_{i}\in\gamma}\sqrt{2\left(1-\mbox{Tr}\left(\rho_{A_{\gamma_{i}}}^{2}\right)\right)},$ (19)
where $\gamma=\left\{\gamma_{i}\right\}$ represents the set of all possible bipartitions $\left\{A_{i}\mid B_{i}\right\}$ of $\left\{1,2,\cdots,n\right\}$.
Now, in our case, for a tripartite pure state $|\psi\rangle\in H_{1}\otimes H_{2}\otimes H_{3}$ there are three possible bipartitions $\gamma=\left\{\left\{1\mid 2,3\right\},\left\{2\mid 1,3\right\},\left\{3\mid 1,2\right\}\right\}$.
Consequently, the gme-concurrence is
$C_{gme}\left({\psi}\right)=\min\left\{\sqrt{2\left(1-\mbox{Tr}\left(\rho_{1}^{2}\right)\right)},\sqrt{2\left(1-\mbox{Tr}\left(\rho_{2}^{2}\right)\right)},\sqrt{2\left(1-\mbox{Tr}\left(\rho_{3}^{2}\right)\right)}\right\}.$ (20)
A.7 Magic
We employ the stabilizer 2-Rényi entropy [39], commonly referred to as the "magic measure". Let $|\psi\rangle$ denote a quantum pure state. Given a specific operator $C$, repeated measurements allow for the estimation of the probability $P\left(s\mid C\right)=\left|\langle s|C|\psi\rangle\right|^{2}$. We define the vector $\vec{s}=\left(\vec{s}_{1},\vec{s}_{2},\vec{s}_{3},\vec{s}_{4}\right)$ comprising four $n$-bit strings, with the binary sum denoted as $\left|\vec{s}\right|\equiv\boldsymbol{s}_{1}\oplus\boldsymbol{s}_{2}\oplus\boldsymbol{s}_{3}\oplus\boldsymbol{s}_{4}$. In this framework, the stabilizer $2-$Rényi entropy is characterized by:
$M_{2}\left(|\psi\rangle\right)=-\log\sum_{\vec{s}}\left(-2\right)^{-\left\|\vec{s}\right\|}Q\left(\vec{s}\right)-\log d,$ (21)
where $Q\left(\vec{s}\right)=E_{C}P\left(\boldsymbol{s}_{1}|C\right)P\left(\boldsymbol{s}_{2}|C\right)P\left(\boldsymbol{s}_{3}|C\right)P\left(\boldsymbol{s}_{4}|C\right)$ represents the expectation value over randomized measurements of the Clifford operator $C$, and $d\equiv 2^{n}$ denotes the dimension of the Hilbert space corresponding to $n-$qubits.